What must be the temperature in Fahrenheit so that it will be twice its value in celsiuCe

Grade:8

1 Answers

Arun
5 years ago
asic conversion formula: F = 9/5C + 32
Our problem: 2C = F

Substitute F in the first formula with 2C from the second, which gives you:
2C = 9/5C + 32

Subtract 9/5C from both sides, which gives:
2C - 9/5C = 32

Simplify the left side (2C = 10/5C): 10/5C - 9/5C = 32, which equals 1/5C = 32

Multiply 5 to each side, which leaves C = 160

So, 160 degrees Celsius will equal 320 degrees Fahrenheit.
 
Regards
Arun